if the difference of two numbers is 18 and if the sum of two numbers is 34, what are the two numbers?
Answer by Earlsdon(6287)

Let the two numbers be a and b.
a-b = 18 Rewrite as: a = b+18 and substitute below:
a+b = 34
(b+18)+b = 34 Simplify.
2b+18 = 34 Subtract 18.
2b = 16 Divide by 2.
b = 8 and a = 8+18 = 26.
The two numbers are: 26 and 8